Telescope and AGw Commissioning at LFBG with IRTC - March 31 - April 1 2009 MST

Observers: J. Hill (mtn), A. Rakich (Tucson), J. Storm (Potsdam)
Telescope Operator: A. Ceranski
Telescope Support: J. Little, G. Holmberg, B. Curtis, T. Shih

Summary

Lost first hour to mounting IRTC. Lost second hours to bypassing jammed stow pin.

The had a reasonably productive time of testing the new software build. (B23) Most things worked. Torsten fixed a couple minor things in GCS. We didn't finish all of the detailed SW tests.

Spent the second half working of the recalibration of the WFS. Hotspot is in the newly advertised position. We measured the pupil position on-axis, but then got confused about how this relates to the exact parameters for GCS.

Found that the nominal probe center is off by ~4-4.5 arcsec from the rotator center. We don't understand why it should be this much (were expecting 1 mm). This really messes up the transformations.

Lost the last 2 hours to clouds coming into the north.

Details

17:30 Bogie problems resolved by disabling the B-right drive. But IRTC still not mounted -- decided to mount IRTC and spend the first hour of observing time.

19:30 IRTC is mounted.

20:00 It (using the first hour) would have been a successful strategy except that the right stow pin got jammed (overcurrent), so we couldn't extend it for changing swing arms and balance. We had to swap wires on the stow pin limit switches to trick the interlocks. That took an extra hour.

20:34 authorize IRTC (from IRAF) and the corresponding irc command producing "Exit 67". Nobody knows what that means, but it went away after Chris and I fiddled with it for a while.

20:45 Had to reboot the IRTC windows computer to get it to connect to the camera. Otherwise, IRTC is working OK. We don't yet know if it is mounted in the correct orientation on the telescope - could be off by 180 degrees. (Remember to make a mark before we take it off.) Window is still off by 90deg - we didn't do anything to it.

21:10 Open dome, and reset encoders. T=-1 degC D=-22 degC, Wind =10-12 m/sec. Photometric.

21:15 pointto M5_0358 ACQUIRE PAM- gs=0 - didn't find

Log file will be written to: LOGs/offset_042046.log

   #############################################################
   ###  Sending the following command sequence to the IRTC:  ###
   #############################################################
   irc OffsetPointing 0. 0. 0. RADEC MOUNT true ABS left
ERROR: Cannot coerce string `' to int
  "res = int(substr(resmsg,8,8))"
     line 138: IRTC$offset.cl
     called as: `offset (off1=0., off2=0.)'

Dave T. gave us a corrected copy of offset.cl that works OK with verbose-mode 0.

21:31 - Reauthorize IRTC to get M3 into position, sending the preset didn't do it as I was expecting. Still a bug for Paul.

21:32 pointto M5_0358 ACQUIRE PAM- gs=0 again. - found star in upper right corner of Acquire image! Adjust *IE= CA= *

21:45 pointto M5_0358 ACTIVE PA=INDEF gs=4 - didn't find off-axis star, even though on-axis is clearly on IRTC.

21:47 pointto M5_0358 ACTIVE PA=INDEF gs=1 - didn't find off-axis star, even though on-axis is clearly on IRTC.

21:50 pointto M5_0358 ACTIVE PA=INDEF gs=0 - locked guiding, but struggling with WFS on this really bright star.

21:55 pointto BS9139 ACTIVE PA=INDEF gs=0 - Telescope stops for bad polys..........

22:00 pointto BS9139 ACTIVE PA=INDEF gs=0 - succeeds, but a lot of trouble with the acquisition offset. Centroid cross is at the wrong place in the acquisition image display. Torsten thinks this is a problem when sextractor finds multiple stars. We apparently still don't have sextractor tuning that can handle defocussed/comatic images.

Some rough collimation gets us to a 0.9 arcsec image.

22:11 pointto BS9139 ACTIVE PA=INDEF gs=1 - couldn't move probe

22:12 pointto BS9139 ACTIVE PA=INDEF gs=2 - couldn't move probe

22:13 pointto BS9139 ACTIVE PA=INDEF gs=3 - Got this off-axis star, and dropped it into the 1.08 arcsec circle on the acquisition offset, and overshot about 0.5 arcsec, but that may have been seeing as the image is very jumpy. GCS cycles until stable is set to 1.

22:21 PSF/PMC hits actuator force limits because of accumulated Z22. Aaron did not have primary mirror ventilation running! Clear C00 and star collimating afresh.

21:38 set hotspot position to Jesper's new lab value of 282.9, 279.6, but that was in AIP_L.cfg, and we are really using AIP_IRTC_L.cfg.

22:40 pointto BS9139 ACTIVE PA=INDEF gs=3 - claimed star was too dim because of some residual aberration blurring it out. Cleared C00 again.

22:43 pointto BS9139 ACQUIRE - Adjust IE=-5 CA=-42

22:46 pointto BS9139 ACTIVE PA=INDEF gs=3

22:55 set hotspot position to Jesper's new lab value of 282.9, 279.6 in AIP_IRTC_L.cfg.

23:08 set edge coordinates in AIP_IRTC_L.cfg to 395.2, 241.1, but not positive I have this exactly correct. Half a lenslet from the center in each direction. But that was wrong because this was an off-axis star. I had it wrong, but Doug will straighten me out later.

23:18 Torsten kills GCS for a patch, but the new build of the patch failed. So back to the previous version.

23:28 pointto BS9148 ACTIVE PA=INDEF gs=0 (on-axis) - succeeded, but pointing off some

/Repository/GCSfiles/dailyimages/wfscimage000088_irc.zern
    ...Done: Filename = /Repository/GCSfiles/dailyimages/wfscimage000089.fits
=> Reading Info file....
=> Rewriting fits file with DD information....
=> Processing file wfscimage000089_dd.fits
% READFITS: EOF encountered attempting to read extension 0
% SXPAR: FITS Header (first parameter) must be a string array
% Error occurred at: SXPAR             130
  /home/lbcobs/LBCFPIA/lbcfpia/lib/astron/pro/sxpar.pro
%                    RECON_PROCESS_IMAGE   59
  /home/LBTO/agw/idl/recon/recon_process_image.pro
%                    RECON_DETERMINE_ZERNIKES  101
  /home/LBTO/agw/idl/recon/recon_determine_zernikes.pro
%                    AGW_POST_COLLIMATE   84
  /home/LBTO/agw/idl/agw/agw_post_collimate.pro
%                    GCS_WFSC_DISPLAY   63
  /home/LBTO/agw/idl/gcs/gcs_wfsc_display.pro
%                    $MAIN$
% Execution halted at: RECON_PROCESS_IMAGE   59
  /home/LBTO/agw/idl/recon/recon_process_image.pro

Maybe this was because Andrew and I were both running the program at the same time.

23:39 guidecam_lowcounts from 1000 to 500. guidecam_highcounts from 45000 to 25000.

Move to BS9148

23:40 pointto BS9148 ACTIVE PA=INDEF gs=0 (on-axis)

23:46 Torsten kills GCS to install a patched version. Among other things it prints Z22 correctly on the GUI.

23:47 pointto BS9148 ACTIVE PA=INDEF gs=0 (on-axis) - didn't find star pointing must have drifted off.

23:50 pointto M5_1332 ACQUIRE PAM- gs=0 (Azimuth unwraps to 117 deg) - IE=-38 CA=-52

23:55 J. Little reports that #1 Chiller in utility building has quit. So we drop back to only left mirror ventilation as we aren't using the right.

00:00 pointto BS9148 ACTIVE PA=INDEF gs=0 (on-axis) - telescope collides with building on slew. HBS wouldn't restart after I reset the MCR relay. Error on HBS was "E22 H#4 not selected". HBS was willing to start in Local, but not after 6 tries in remote. Aaron stops and restarts MCSPU and then HBS is willing to start. Lost 20 minutes.

00:15 John L. and Aaron got the chiller working again. We drop to only ventilating the SX mirror to lighten the load.

Checking the new hotspot position on the guide camera

00:20 pointto BS9148 ACTIVE PA=INDEF gs=0 (on-axis) ProbeXY -0.587,+611.844 for on-axis

Seeing is 00:38 +1 mm Y = 612.844

00:40 -1 mm Y = 610.844

00:42 Block guiding in PCS

00:43 +2 mm Y - 612.844 - no light as expected

00:45 +1.5 mm Y 612.344

00:47 +1 mm Y 611.844

Restart collimating with IDL, and restart guiding with probe at -0.587 610.844

00:55 Primary having trouble getting to position, pressed collimate to update it, but then got a pile of correction. We puzzled for a minute, but this is because we had gone 20 minutes without any secondary offloads.

01:03 Blocked guiding

01:04 +1 mm Y = 611.844

01:06 -1 mm Y = 609.844

01:09 Restart guiding and collimating at 610.844

01:11 Blocked guiding

01:12 -1 mm X = -1.587

01:13 +1 mm X +0.413

01:14 restart guiding at -0.587 610.84

Some elongation in wfsimage 0021

1:18 Blocked guiding

1:19 +1 mm Y = 611.844

1:20 -1 mm Y = 609.844

Conclude that we are centered as well as we have to be, although the measurement is not especially precise in this seeing. We certainly agree with Jesper's revised hotspot position of 282.9, 279.6 .

01:22 pointto BS9148 ACTIVE PA=INDEF gs=0 (on-axis) - didn't find a star

Move to BS9152

01:25 pointto M5_1332 ACQUIRE PAM- gs=0 - Adjust IE=-13 CA=-59

01:32 pointto BS9152 ACTIVE PA=INDEF gs=0 (on-axis)

01:42 pointto BS9152 GUIDE gs=17, - no star

01:43 pointto BS9152 GUIDE gs=18, - no star

01:42 pointto BS9152 GUIDE gs=2 - found star

01:43 pointto BS9152 GUIDE gs=1 - found star if slightly different place

01:51 Stop and restart PCS to pickup the LEFTZEROPOINT of -187.703 copied from LUCIFER (My typo was -187.723.)

01:52 pointto BS9152 GUIDE gs=1 - found star

Seeing has degraded to 2 arcsec, or maybe Andrew hasn't focussed in a while (both).

We can't seem to manage a connection to Jesper via Polycom or Skype.

02:03 pointto BS9145 ACTIVE gs=0

02:12 pointto BS9145 GUIDE gs=7 - found star far off

02:15 pointto BS9145 ACQUIRE gs=0

02:16 offsetting N-S moves the star reliably N-S on the acquisition image, so the AGw is bolted on correctly.

02:20 Jesper finally gives up and connects via the conventional telephone. He is checking to make sure we still have the "good" transformation in oacserver.

02:21 Adjust *IE=-18 CA=-65 *

02:25 pointto BS9145 GUIDE gs=8 - found star off by 6 arsec - probe was at -1.63 515.10

02:30 Jesper thought we might have a wrong X parameter in the configuration, so restart oacserver, GCSL. Home the AGw for good measure.

02:33 pointto BS9145 GUIDE gs=0 - OK

02:35 pointto BS9145 GUIDE PA=INDEF gs=7 - found star but almost at the left edge of the image

02:38 pointto BS9145 GUIDE PA=0 gs=7 - couldn't move probe there

02:39 pointto BS9145 GUIDE PA=180 gs=7 - didn't find star

02:45 Authorize as LUCIFER instead of IRTC

02:46 pointto BS9145 GUIDE PA=180 gs=0

02:50 cl < home$spiral.cl found at offset 0 -30. Adjust IE=-28 CA=-22 for LUCIFER pointing model.

02:58 pointto BS9145 GUIDE PA=180 gs=0 (IRAF gets all hosed up)

03:01 pointto BS9145 ACTIVE PA=180 gs=0

03:05 pointto BS9145 ACTIVE PA=180 gs=7 - no star on acq image

03:07 pointto BS9145 ACTIVE PA=INDEF=0 gs=0 - drifted? or did I send the wrong star

03:10 ACQUIRE at PA=180 - don't find star

03:13 ACQUIRE at PA=0 - Adjust IE=-35 CA=-20

03:16 ACQUIRE at PA=180 - star moves in ~X by ~9 arcsec; So probe is not pointing at the rotator center!

03:20 ACTIVE at PA=0 gs=0 - Chase failed to make tolerance

03:34 Stopped IDL collimation and restarted GCS WFS

03:42 Lost star when we hit collimation limit Y=-2.6 (did GCS diverge?) The issue was a bank of clouds coming up from the north.

03:57 pointto BS9166 ACTIVE gs=0

04:00 HBS stops and we hit the building again We will close for clouds as soon as Aaron gets the oil going again.

04:10 Close for clouds........

05:34 Tried to take IRTC picture, but there was a problem with the socket connection.

Apr  1 05:44:44 lbtmu05 irc[32756]: GetFilter result status: ERROR [Error] - IRSocket::IRSocket -> [::connect] Could not bind to port.

The very short plan for tomorrow

1) Finish the software tests.

2) Verify the parameters for the pupil position in GCS

3) Adjust the AGw probe encoder zeropoints to get probe on the rotator center.

See Joar's document for the long plan.........

-- JohnHill - 01 Apr 2009
Topic revision: r5 - 02 Apr 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback